Rajdhini express moves at a speed of 150 km/h. How long will it take to cover a distance of 15 km?

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
Dear student

Sᴏʟᴜᴛɪᴏɴ :-

 

We know that,

 

 Distance = Speed × Time

 

D = s × t

 

Put the above given values in it, we get,

 

 15 = 120 × t

 

 t = 15/120

 

 t = 1/8 hr

 

We know that,

 

 1 hour = 3600 s

 

So,

 

 1/8 hrs = 1/8 × 3600

 

 3600/8

 

 450 s

 

Hence,

 

  • Time taken (t) = 450 seconds or 7.5 min